OpenStudy (anonymous):

Let f(x) = 13 x^3 - 10 x^2 Find all values of x (if any) for which f may have a relative maximum or a relative minimum.

OpenStudy (cwrw238):

differentiate and equate to 0 solve for x to find x coods of turning points identify the nature of turning points

OpenStudy (anonymous):

turning points?

OpenStudy (anonymous):

i dont think we're that far yet

OpenStudy (cwrw238):

yea - max, minm or points of inflection on the graph of the function - you are learning calculus right?

OpenStudy (anonymous):

oh ok and yes calculus

OpenStudy (cwrw238):

f'(x) = 39x^2 - 20x = 0 x(39x - 20) = 0 x = 0 or 20/39

OpenStudy (cwrw238):

so there is a t.p. at x=0 and x = 20/39 to find their nature find f" can you continue from here?

OpenStudy (anonymous):

find the 2nd deriv from the first deriv equation?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

yeah i can go from here but im doing my homework and they are asking me for something odd like . g'(x)>0 for the following values of x: and its multiple choice like a<x<b

OpenStudy (cwrw238):

oh -maybe they are asking you to find maxm or minm by the sign of the derivative either side of the two values of x for example if at x = 0 f'(x) is >0 just before that point and < 0 just after then x=0 is a maximum

OpenStudy (anonymous):

well ive found the critical points but i dont know how to implement them to find what they're looking for

OpenStudy (cwrw238):

can you post the original question?

OpenStudy (anonymous):

here it is : Let f(x) = 13 x^3 - 10 x^2 Find all values of x (if any) for which f may have a relative maximum or a relative minimum. Give answers in increasing order (smallest first) there should be two answers

OpenStudy (cwrw238):

ok - we've done that. they are x = 0 and x = 20/39 They haven't asked us to identify which is max or min though we can do that if we like.

OpenStudy (cwrw238):

f"(x) = 78x - 20 f"(0) is < 0 so this is a maxm f(20/39) is >0 so this minm.

OpenStudy (anonymous):

yeah i plugged in x=0 and x = 20/39 and it says its right

OpenStudy (anonymous):

thank you very much, you saved my life lol

OpenStudy (cwrw238):

yw